### v2

# Search for Solar Boosted Dark Matter Particles at the PandaX-4T Experiment Guofang Shen

Search for Solar Boosted Dark Matter Particles at the PandaX-4T Experiment
Guofang Shen
School of Physics, Beihang University, Beijing 102206, China
Zihao Bo
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Wei Chen
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xun Chen
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yunhua Chen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zhaokan Cheng
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Xiangyi Cui
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Yingjie Fan
Department of Physics, Yantai University, Yantai 264005, China
Deqing Fang
Key Laboratory of Nuclear Physics and Ion-beam Application (MoE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Zhixing Gao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Lisheng Geng
lisheng.geng@buaa.edu.cnSchool of Physics, Beihang University, Beijing 102206, China
Peng Huanwu Collaborative Center for Research and Education, Beihang University, Beijing 100191, China
International Research Center for Nuclei and Particles in the Cosmos & Beijing Key Laboratory of Advanced Nuclear Materials and Physics, Beihang University, Beijing 100191, China
Southern Center for Nuclear-Science Theory (SCNT), Institute of Modern Physics, Chinese Academy of Sciences, Huizhou 516000, China
Karl Giboni
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xunan Guo
School of Physics, Beihang University, Beijing 102206, China
Xuyuan Guo
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zichao Guo
School of Physics, Beihang University, Beijing 102206, China
Chencheng Han
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Ke Han
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Changda He
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinrong He
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Di Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Houqi Huang
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Junting Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Ruquan Hou
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yu Hou
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xiangdong Ji
Department of Physics, University of Maryland, College Park, Maryland 20742, USA
Xiangpan Ji
School of Physics, Nankai University, Tianjin 300071, China
Yonglin Ju
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Chenxiang Li
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jiafu Li
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Mingchuan Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Shuaijie Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Tao Li
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Zhiyuan Li
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Qing Lin
State Key Laboratory of Particle Detection and Electronics, University of Science and Technology of China, Hefei 230026, China
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Jianglai Liu
jianglai.liu@sjtu.edu.cnNew Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Congcong Lu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xiaoying Lu
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Lingyin Luo
School of Physics, Peking University, Beijing 100871, China
Yunyang Luo
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Wenbo Ma
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Yugang Ma
Key Laboratory of Nuclear Physics and Ion-beam Application (MoE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Yajun Mao
School of Physics, Peking University, Beijing 100871, China
Yue Meng
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xuyang Ning
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Binyu Pang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Ningchun Qi
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zhicheng Qian
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiangxiang Ren
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Dong Shan
School of Physics, Nankai University, Tianjin 300071, China
Xiaofeng Shang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiyuan Shao
School of Physics, Nankai University, Tianjin 300071, China
Manbin Shen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Wenliang Sun
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yi Tao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Anqing Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Guanbo Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Hao Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jiamin Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Lei Wang
College of Nuclear Technology and Automation Engineering, Chengdu University of Technology, Chengdu 610059, China
Meng Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Qiuhong Wang
Key Laboratory of Nuclear Physics and Ion-beam Application (MoE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Shaobo Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Siguang Wang
School of Physics, Peking University, Beijing 100871, China
Wei Wang
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Xiuli Wang
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xu Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Zhou Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yuehuan Wei
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Weihao Wu
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yuan Wu
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Mengjiao Xiao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiang Xiao
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Kaizhi Xiong
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yifan Xu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Shunyu Yao
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Binbin Yan
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Xiyu Yan
School of Physics and Astronomy, Sun Yat-Sen University, Zhuhai 519082, China
Yong Yang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Peihua Ye
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Chunxu Yu
School of Physics, Nankai University, Tianjin 300071, China
Ying Yuan
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zhe Yuan
Key Laboratory of Nuclear Physics and Ion-beam Application (MoE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Youhui Yun
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xinning Zeng
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Minzhen Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Peng Zhang
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Shibo Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Shu Zhang
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Tao Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Wei Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Yang Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Yingxin Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Yuanyuan Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Li Zhao
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Jifang Zhou
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Jiaxu Zhou
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Jiayi Zhou
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Ning Zhou
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xiaopeng Zhou
xpzhou@buaa.edu.cnSchool of Physics, Beihang University, Beijing 102206, China
Yubo Zhou
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zhizhen Zhou
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Haipeng An
anhp@mail.tsinghua.edu.cnDepartment of Physics, Tsinghua University, Beijing 100084, China
Key Laboratory of Particle and Radiation Imaging (MoE), Tsinghua University, Beijing 100084, China
Haoming Nie
Department of Physics, Tsinghua University, Beijing 100084, China
(May 13, 2025)
Abstract
We present a novel constraint on light dark matter utilizing tonneyear of data acquired from the PandaX-4T dual-phase xenon time projection chamber. This constraint is derived through detecting electronic recoil signals resulting from the interaction with solar-enhanced dark matter flux. Low-mass dark matter particles, lighter than a few MeV/, can scatter with the thermal electrons in the Sun. Consequently, with higher kinetic energy, the boosted dark matter component becomes detectable via contact scattering with xenon electrons, resulting in a few keV energy deposition that exceeds the threshold of PandaX-4T. We calculate the expected recoil energy in PandaX-4T considering the Sun’s acceleration with heavy mediators and the detection capabilities of the xenon detector. The first experimental search results using the xenon detector yield the most stringent upper limits cross-section of at / for a solar boosted dark matter mass ranging from to /, achieving a 23 fold improvement compared with earlier experimental studies.
Although astrophysics and cosmology studies have accumulated ample evidence that much of the universe’s mass is dark matter (DM), its nature remains largely obscure Bertone et al. (2005); Bertone and Hooper (2018); Aghanim et al. (2020).
The freeze-out mechanism proposes that dark matter was formed when particles frequently collided within a dense, high-temperature ”soup” in the early universe. In this model, most particles initially exist as dark matter but later are destroyed. The remaining DM particles are usually classified as weakly interacting massive particles (WIMPs) Steigman et al. (2012).
In recent decades, WIMPs have emerged as the predominant candidates for elucidating the phenomenon of missing mass, notwithstanding the lack of compelling evidence from direct detection experiments utilizing xenon or other materials Undagoitia and Rauch (2015); Liu et al. (2017); Roszkowski et al. (2018); Schumann (2019). In recent years, xenon-based detectors such as PandaX-4T, LZ, and XENONnT have provided the most stringent limits on WIMPs–nucleon spin-independent interactions, pushing the cross-section down to for a DM mass of , for , and for , respectively Bo et al. (2024); Aalbers et al. (2024); Aprile et al. (2023).
These results present significant challenges to the mainstream supersymmetry framework Bagnaschi et al. (2015) in the WIMP sector. However, there remains considerable potential for further exploration in lower-mass regions.
The traditional method for directly detecting DM is to measure the elastic scattering between target nuclei and DM particles. Since cold DM in the halo is non-relativistic, the direct detection strategy is only sensitive to DM with masses exceeding sub-GeV due to the detector thresholds. In recent years, pushing the search to lower mass scales of sub-GeV or even sub-MeV DM, which are still untapped for promising thermal relic DM possibilities, has spurred many studies. Various approaches have been proposed, such as accessing the electron recoil (ER) channel Essig et al. (2012a); Emken et al. (2019); Essig et al. (2012b, 2016); Agnes et al. (2018a); Amaral et al. (2020); Barak et al. (2020); Aguilar-Arevalo et al. (2019); Cheng et al. (2021a), lowering the nuclear recoil (NR) detection threshold Pirro and Mauskopf (2017); Abdelhameed et al. (2019); Mancuso et al. (2020); Aprile et al. (2019a); Agnes et al. (2018b), and utilizing the so-called Migdal effect Ibe et al. (2018); Baxter et al. (2020); Essig et al. (2020); Akerib et al. (2019); Armengaud et al. (2019); Liu et al. (2019); Aprile et al. (2019b).
Currently, considerable attention has been paid to specific processes that would boost the kinetic energy of DM particles Necib et al. (2017); Kouvaris (2015a); Cui et al. (2022); Aoki and Toma (2018); Giudice et al. (2018); Kachulis et al. (2018); Fornal et al. (2020); Wang et al. (2020); Bondarenko et al. (2020); Calabrese et al. (2022); Ge et al. (2021). For example, the search for DM boosted by cosmic rays presents constraints on DM-electron cross-sections () for the unexplored DM mass from to in PandaX-4T Shang et al. (2024).
Another novel process called solar boosted dark matter (SBDM) scenario has been proposed An et al. (2018); Emken et al. (2018); An et al. (2021) to open the window for sub-MeV DM detection. The Sun can act as an accelerator of DM particles via evaporation or scatterings with the thermal plasma in the core of the Sun where the temperature could reach keV Kouvaris (2015b); Emken et al. (2018); An et al. (2018, 2021).
In this letter, we focus exclusively on the generation and detection of SBDM through electron scattering An et al. (2018, 2021), as the contribution of ions to MeV scale SBDM is negligible Emken (2022).
The ambient DM with a typical velocity between and would be attracted by the Sun via gravity when they pass nearby. The tremendous amount of energetic electrons in the Sun with a density at the level of could effectively heat the DM particles and boost their kinetic energy to keV levels An et al. (2021). The SBDM could then escape the Sun and reach terrestrial detectors like PandaX-4T Meng et al. (2021) leaving detectable signals that exceed the threshold with the same amplitude on the bounded electrons of the xenon. Calculating the flux of SBDM emanating from the surface of the Sun is imperative. Subsequently, we will assess the event rate in the xenon detector based on the flux and for different DM masses.
The SBDM flux arriving at the earth can be estimated as An et al. (2018)
where is the DM flux in the local Milky Way halo with a DM local density Catena and Ullio (2010) and a typical DM velocity distribution, and is the scattering radius determined through the radius-temperature relation Bahcall et al. (2005). indicates the gravitational focusing effect, which enhances the solar scatterings and has a value of O() when An et al. (2018). is the Sun-Earth distance. The is the electron number density within the radius of the Sun’s core (). The mean free path for a DM particle inside the core of the Sun is which would dramatically affect the calculations of the flux when .
The possibility of multiple scattering should be considered in Eq. 1 but is difficult to treat analytically. So, a Monte Carlo simulation program An et al. (2018) is built to mimic the propagation of DM particles inside the Sun with an elaborate solar model taken from Serenelli et al. (2009).
In the simulation, the perpendicular distance of halo DM particles to the Sun, which is known as impact parameter , is confined within to incorporate gravitational focusing in the calculation of the SBDM. Beyond this range, only a small proportion of the slowest DM could be trapped by the gravity field, minimally contributing to the boosted flux. is the simulated kinetic energy distribution of an SBDM particle, where , the DM mass, and are two key input parameters for the simulation. The boosted flux observed at the detector’s location is illustrated as
where is the area where the DM particles could be boosted in the simulation and .
Taking into account , and , the SBDM spectra, with = and , are shown in Fig. 1. The kinetic energy transfer efficiency reaches a maximum when the masses of the elastically scattering particles are equal. The spectra for
= 0.5 MeV are harder than those for the other three DM masses.
Figure 1: SBDM kinetic energy spectrum. The first value in parentheses in the legend above represents the dark matter mass in units of MeV, and the value adjacent is in units of .
According to the DM-electron scattering analysis Essig et al. (2012b), the velocity-averaged electron recoil differential cross-section for ionization from an initial atomic state
with principal and angular quantum numbers and can be expressed as
where is the DM-electron reduced mass, and is the recoil energy. The DM form factor can take different values depending on the momentum transfer . The atomic form factor,
, describes the strength of the ionization process from the atomic state (). In this work, we consider An et al. (2021), which means that the interaction between SBDM particles and electrons is contact. is the flux-average of squared inverse speed for the SBDM with a velocity larger than that could generate a minimum recoil.
There is a framework Essig et al. (2016) built to calculate sub-GeV dark matter scattering off electrons with
semiconductor targets.
Our entire calculation is based on the framework to calculate the DM-electron scattering rate, but with the following modifications.
First, we use xenon as the target; second, we sum the electron recoiling and deexcitation photon energy from different electron shells in xenon; last,
the atomic form factor is taken from Ref. Catena et al. (2020), also used in cosmic-ray boosted DM research in PandaX Shang et al. (2024).
Next, we calculate the differential recoil rate
where is the electron number density in the xenon target. is the differential scattering cross-section, and the is the SBDM flux at the detector’s location An et al. (2018).
PandaX-4T Meng et al. (2021) is located in the China Jinping Underground Laboratory (CJPL) Kang et al. (2010); Guo et al. (2021). The central detector is a dual-phase time projection chamber (TPC) with tonnes of liquid xenon in the sensitive region. The detector is defined by reflective polytetrafluoroethylene (PTFE) panels, separating mm between opposing panels. The internal electric fields are established by a cathode grid at the bottom, along with a gate and anode mesh positioned just below and above the liquid level, maintaining vertical spacings of and mm, respectively. In addition, two arrays of and three-inch Hamamatsu R11410-23 photomultiplier tubes (PMT) and are installed on the top and bottom of the TPC.
Electron recoil energy is converted into prompt scintillation () in liquid xenon and delayed proportional electroluminescence photons () in gaseous xenon. The signals and are collected by the PMT arrays and used to reconstruct the energy and three-dimensional position. To enhance our understanding of the detector response, we utilize two primary calibration methods: the calibration of electronic and nuclear recoil events. The ER response is calibrated by the 220Rn Ma et al. (2020) and 83mKr Zhang et al. (2022) and other radioactive sources such as 60Co and 137Cs. The Am-Be neutron source and the Deuterium-Deuterium neutron generator are used for NR calibration. More details are given in Bo et al. (2024).
The electron equivalent energy of each event in the detector is reconstructed as
where is the average work function of liquid xenon, with a value of 13.7 eV from Szydagis et al. (2011). The is the bottom-only . The use of is to avoid the effects of
saturation and dead channels of the top PMTs.
PDE, EEE, and represent the photon detection efficiency, electron extraction efficiency, and single electron gain, respectively. The here already involve the PDE for S2 and are obtained by measuring the smallest recorded signals Meng et al. (2021). The values of PDE and EEE are predetermined by calibrating the monoenergetic electronic recoil peaks: 41.5 keV (83m Kr), 164 keV (131mXe), 236 keV (129mXe), and 408 keV (127Xe).
Fig. 2 illustrates the expected interaction rate for different DM masses with representative cross-sections and the total efficiencies in PandaX-4T. The total efficiencies are estimated using both the data-driven and waveform-simulation method Li et al. (2024); Luo et al. (2024). Furthermore, the energy resolution curves of Run0 and Run1 are given in Fig. 2. The structure of the energy spectrum is attributed to interactions between the SBDM and different shell electrons of the xenon. The recoil energy could reach four keV, beyond the PandaX-4T experiment threshold Meng et al. (2021); Bo et al. (2024).
Figure 2: The expected interaction rate of SBDM in the xenon detector with = 0.25 MeV, 0.5 MeV, and 1.0 MeV for . For = 0.5 MeV, the expected interaction rate with from to cross-sections is also shown. The dashed green and dashed orange lines represent the total efficiencies of Run0 and Run1, and the shield bands illustrate the efficiency uncertainties. The dotted and dashed magenta curves are the energy resolution of Run0 and Run1, respectively. With interaction rate, efficiency and energy resolution convolved, the corresponding reconstructed signal spectrum could be derived and used in the following analysis.
This analysis uses the same data sets from Run0 and Run1 as Ref. Bo et al. (2024), including 11 subsets segmented according to slightly different electric field configurations and background levels. The fiducial masses are tonne for Run0 and tonne for Run1. The total exposure is tonneyear.
In the search for SBDM, the signals are expected to manifest themselves as electronic recoil events, in contrast to conventional WIMP signals. Consequently, an accurate estimation and characterization of all ER backgrounds is essential for the statistical analysis and interpretation of SBDM signals.
The energy of SBDM signals ranges from to keV. To enhance the measurement of the ER background from 222Rn, we choose the final region-of-interest (ROI) from 0 to 30 keV. Furthermore, a threshold of 99.5% ER acceptance cut was implemented to exclude most NR events arising from solar 8B neutrinos and neutron interactions with the xenon nucleus as well as radon plateout on the surface of the PTFE reflector of the PandaX-4T TPC. This cut is derived from the simulated data of the PandaX-4T signal model Luo et al. (2024). Finally, we consider the primary ER backgrounds, including tritium, 85Kr, 214Pb, 212Pb, 127Xe, 136Xe, 124Xe, neutrino, material, and accidental events. A comprehensive examination of each component is presented in Zeng et al. (2024).
Run
Run0
Run1
Exposure
198.9 tonneday
363.3 tonneday
Expected
Fitted
Expected
Fitted
Tritium
575.1 32.8
214Pb
327.2 18.8
328.1 17.1
212Pb
57.8 14.7
85Kr
94.2 47.3
Material
49.4 3.3
136Xe
36.9 2.5
127Xe
6.1 0.3
124Xe
2.3 0.6
Solar
43.0 4.6
Accidental
7.6 2.4
Signal
Total fitted
1196.5 32.7
1439.4 36.2
Observed
1197
1431
Table 1: Summary of the expected background and signal, and the best fit of the background plus signal hypothesis for each component in Run0 and Run1. Systematic uncertainties associated with the signal model and detection efficiencies are incorporated into the reported uncertainties.
Data from the PandaX-4T experiments, specifically the
1197 (Run0) and 1431 (Run1) electron recoil (ER) events, have been meticulously selected for this analysis. A profile likelihood ratio (PLR) approach Baxter (2021) is used in this analysis with a one-dimensional signal energy spectrum based on HistFitter Baak et al. (2015). In this analysis, both background-only and background-plus-signal hypotheses are tested. Background contributions are well aligned with the expected values, and all nuisance parameters remain within of the input values. The best fit of the SBDM result is illustrated in Fig. 3 and the expected events from each component are summarized in Table 1. Ultimately, no significant SBDM signal is detected above the background.
Figure 3: Best combined fit results of Run0 and Run1 data with background-plus-signal hypothesis. Error bars are statistical uncertainties.
The 90% confidence level (C.L.) two-side upper limit exclusion (solid red line) of the SBDM parameter space, along with the sensitivity (green band), is presented in Fig. 4.
For comparison, recent results from the phenomenological study An et al. (2018), the SBDM analysis utilizing the HPGe detector from CDEX Zhang et al. (2024), and the constraints derived from cosmological and astrophysical observables related to freeze-out DM Planck Collaboration et al. (2016), and other analyses of sub-GeV DM Li et al. (2023); Cheng et al. (2021b); Aprile et al. (2019c); Essig et al. (2012b) are overlaid. Our results cover a large region from 20 keV/
to 10 MeV/, with cross-sections ranging between and
. The lowest 90% confidence level on the SBDM cross-section is for an SBDM mass of . Our results provide the most stringent constraints, achieving a 23-fold improvement compared to recent experimental studies conducted by CDEX Zhang et al. (2024).
Figure 4: 90% confidence level (C.L.) exclusion upper limits (red line) and exclusion region (lightgreen region) for the DM-electron cross-section are presented for SBDM. The light gray band is stellar cooling constraints from red giant (RG) stars Chang et al. (2021). The gray curve denotes the freeze-out DM mechanism Planck Collaboration et al. (2016), while the dashed yellow line indicates the SBDM limits from Ref. An et al. (2018). Additionally, the orange line is from the CDEX experimental result Zhang et al. (2024), and the direct constraints based on S2 analysis from Ref. Li et al. (2023); Cheng et al. (2021b); Aprile et al. (2019c); Essig et al. (2012b) are also overlaid at the right side.
In summary, we have performed the direct detection of DM-electron scattering by searching for an energetic
SBDM flux produced through rescattering in the
Sun with the 1.54 tonneyear data of Run0 and Run1 from PandaX-4T. This leads to a new sensitivity at the sub-pb level for light DM in the sub-MeV mass range. No significant dark matter signals are identified above the expected background. A new upper limit exclusion is established for the SBDM-electron interactions with heavy mediators, robustly excluding sub-MeV DM with a scattering cross-section with electrons within to . The well-known freeze-out mechanism below the region could be excluded. Combined with analyses such as ionization signals Li et al. (2023),
The future upgraded PandaX-4T detectors, with significantly enhanced sensitivity and optimized S2-only event detection thresholds Abdukerim et al. (2024), will explore the narrow unconstrained parameter window in the mass range of .
We thank Josef Pradler and Shao-Feng Ge for their helpful discussions. This project is supported in part by grants from National Science Foundation of China (Nos. 12090060, 12090061, 12105008, 12435007, U23B2070), a grant from the Ministry of Science and Technology of China (Nos. 2023YFA1606200, 2023YFA1606201, and by Office of Science and Technology, Shanghai Municipal Government (grant No. 22JC1410100, 21TQ1400218), the National Natural Science Foundation of China under Grant No. 12435007. We thank for the support by the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Hongwen Foundation in Hong Kong, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.
Szydagis et al. (2011)M. Szydagis, N. Barry,
K. Kazkaz, J. Mock, D. Stolp, M. Sweany, M. Tripathi, S. Uvarov, N. Walsh, and M. Woods, Journal of Instrumentation 6, P10002 (2011).